# Long-Tailed 3D Detection via 2D Late Fusion

### v3

@@ Appendix B Transformer-Based LiDAR Detectors
Table 7: Per-class performance on nuScenes. Our late-fusion approach achieves the highest per-class AP on 5 out of 10 classes. Compared to DeepInteraction, our approach improves construction worker by 25.2%, stroller by 6.8%, and pushable-pullable by 27.3%. Note, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and Stro. is stroller. We highlight classes with Medium and Few examples per class in blue.
Method Car Adult Truck CV Bicy. MC Child CW Stro. PP
@@
CP + FCOS3D [5] 88.5 86.6 63.4 29.0 58.5 68.2 5.3 35.8 31.6 39.3
CP + DINO Det. (Ours) 86.3 86.2 60.6 35.3 70.1 75.9 8.8 55.9 37.7 58.1
Appendix D Inference Runtime
Table 6 compares the inference runtime of our method against prior works on a single GPU. Although late-fusion methods such as our method (CP + DINO) and “CP + FCOS3D” [5] run two uni-modal detectors, we run them in parallel. Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [21], TransFusion [31] and DeepInteraction [65]. However, running two detectors in parallel will require more memory usage. Importantly, late-fusion methods including CP + DINO and [5] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck. This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
Appendix E Per-Class Breakdown Results
We compare recent multi-modal methods w.r.t per-class AP in Table 7. All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community. Notably, our late-fusion approach achieves 20% higher AP on pushable-pullable and 6% higher AP on stroller than prior work. In general, our late-fusion approach yields considerable improvement on classes with medium and few examples. Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8% AP, which is still much better than compared methods. We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
Appendix F Analysis of Misclassification
For 3D detection, localization and classification are two important measures of 3D detection performance. In practice, we cannot achieve perfect performance for either. In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detect but misclassify a child as adult versus not detecting this child). Therefore, we evaluate CenterPoint and CMT using the hierarchical AP metric introduced by [5], which considers such semantic relationships across classes to award partial credit. Applying this hierarchical AP reveals that classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses.
Table 8: Diagnosis using the mAPH metric on selected classes. We analyze CenterPoint and CMT with and without 2D late-fusion. Comparing the rows of LCA=0 for CenterPoint and CMT with and without late-fusion, we see our techniques bring significantly improvements on classes with medium and few examples such as construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable (PP). Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, confirming that objects from rare classes are often detected but misclassified as some sibling classes.
Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller PP
CenterPoint (LiDAR-Only) LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
w/ Hierarchy LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
CenterPoint LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
w/ Hier. + DINO Det. LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
  CMT LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
CMT-L LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
+ DINO Det. LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
Table 9: Evaluating 2D detection performance. We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes. However, 2D RGB detectors achieve higher performance. Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane. This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors. All detectors in this table are only trained on the nuScenes train-set.
@@ Appendix F Analysis of Misclassification
Appendix G Analysis of 2D Detection Performance
Table 9 shows that 2D detectors (e.g., DINO) outperform state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set. Importantly, DINO performs significantly better than BEVFormer (15.9 vs. 2.1 mAP) on rare classes.
Appendix H Results on Standard nuScenes Benchmark
We evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [2]. Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [8] and DINO [55] yields a 2.6% and 6.9% performance increase for mAP and NDS on the standard nuScenes benchmark, respectively. Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories. Next We apply our late-fusion to CMT [66], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements. Late-fusion of CMT [66] and DINO [55] only marginally improves bicycle and traffic cone performance. This suggests that CMT is already able to effectively utilize multi-modal information to correctly localize and classify common classes.
Table 10: 3D detection performance on the standard nuScenes benchmark. We evaluate the impact of late-fusion for the standard nuScenes benchmark. Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark. Notably, fusing CenterPoint [8] with DINO [55] improves detection accuracy by 2.6% mAP and 6.9% NDS. However, fusing CMT [66] (a multi-modal 3D detector) with DINO [55] increases detection accuracy by 0.2%, only improving on bicycle and traffic-cone. We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion. Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
@@ Appendix H Results on Standard nuScenes Benchmark
CMT [66] + DINO Det. 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
Table 11: Comparison with the Argoverse 2 state-of-the-art. We present results on AV2 evaluated at 50m. FCOS3D [7] achieves poor performance, likely due to inaccurate depth estimates. In contrast, CenterPoint achieves strong performance on all classes. Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3% improvement averaged over all classes. These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
@@
       CenterPoint + FCOS3D [5]        C + L        48.4        79.0        51.4        35.3
       CenterPoint + DINO Det. (Ours)        C + L        52.3        89.4        54.2        38.7
Appendix I Results on Argoverse 2
We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research. AV2 evaluates on 26 classes, which follow a long-tailed distribution. Following [5], we train on and evaluate detections up to 50m. As shown in Table 11, our main conclusions from nuScenes still hold for AV2. FCOS3D [7] yields poor performance on all classes, likely due to inaccurate depth estimates. CenterPoint performs considerably better, achieving high accuracy on classes with many examples. Notably, CenterPoint performs better on AV2’s rare classes (30.2 AP) compared to nuScenes’s rare classes (3.5 AP), likely because AV2 has more examples per-class in-the-tail. Lastly, our proposed late-fusion approach yields an 8.3% improvement over CenterPoint and a 3.9% improvement over prior work. These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Figure 5: Additional qualitative results. We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO. In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance. Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
Appendix J More Visualizations
We present additional visualizations of our late-fusion approach in Fig. 5. Our method correctly classifies geometrically similar but semantically different categories like adult-vs-stroller, bicycle-vs-personal mobility, child-vs-adult, and adult-vs-construction worker. We provide a video of our results in the supplement.
Figure 6: Examples of failure cases. Both our method (columns 1-3) and TransFusion [31] (columns 4 -5) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, the late-fusion predictions miss these cars because our method throws away unmatched RGB detections since we do not have accurate 3D information. In the third row, we see that although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted categorical label. TransFusion also misclassifies this object as an adult.
Failure cases. We visualize common failure cases of our late-fusion approach and compare them with the failure cases of TransFusion [31], an end-to-end trained multi-modal detector. We find that our method fails in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies the object. See Fig. 6 for detailed analysis.
Figure 7: Impact of 2D detector performance on LT3D. While nuScenes is a 3D detection benchmark, we can generate 2D annotations using the provided sensor extrinsics by projecting the 3D annotations onto the 2D image plane. Results show that evaluating 2D detectors using these 2D annotations is a good proxy task (x-axis) that is positively correlated with the downstream 3D detection performance of the full late-fusion pipeline (y-axis). Concretely, training 2D detectors with more data (e.g., training with nuScenes + nuImages vs. nuScenes only) and using better 2D detectors (e.g., DINO vs. YOLO-v7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our late-fusion algorithm.
Appendix K Limitations and Future Work
Our work focuses on LT3D, a problem that emphasizes 3D object detection for rare classes like stroller and debris, which are often safety-critical for AVs. Therefore, improving LT3D is important for ensuring safe autonomy. However, our work does not directly study how addressing LT3D affects downstream perception tasks. Future work should address this limitation. Moreover, as shown in Fig. 7, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance. We find that 2D detection accuracy on nuScenes is a strong proxy for final 3D LT3D performance. Recent works in large-scale vision language models [69, 70, 71, 72] show promising zero-shot results in detecting rare classes. Identifying ways of incorporating foundation models into our late-fusion framework might improve LT3D further.
In addition, our proposed late-fusion approach is unlikely to work well for objects that are visible in LiDAR but occluded in RGB images (or visible in RGB images but occluded in LiDAR, but this is unlikely because LiDARs are typically mounted higher than RGB cameras). Notably, due to LiDAR and RGB sensors not being co-located on the autonomous vehicle, some objects may be visible in one modality but occluded in another. If an object is visible in LiDAR but not in RGB, we would down-weight the confidence scores of LiDAR-based detections that not supported by RGB-based detections. In contrast, if an object is visible in RGB but not in LiDAR, our 2D late-fusion pipeline would not be able to detect this object in 3D.
Next, we note that the success of late-fusion requires good sensor calibration between the RGB cameras and LiDAR sensor. In fact, sensor calibration errors can also affect early- and mid-fusion methods as they require calibrated extrinsics [22]. As publicly available datasets (e.g., nuScenes and Argoverse 2) are collected in real-world scenarios, one may assume that such datasets already capture realistic calibration errors. Nevertheless, studying how sensor calibration errors affects LT3D performance is worthy of further exploration in future work, which requires collecting new datasets by controlling for such errors.
Lastly, we note that although our proposed approach is highly relevant to 3D perception for autonomous vehicles, our work does not evaluate on real robots. However, following prior literature, evaluating on nuScenes and Argoverse 2 serves as a well established proxy for performance in-the-wild.

### v4

@@ F Analysis of Hierarchical Training
TABLE X: Per-class performance on nuScenes. Our late-fusion approach achieves the highest per-class AP on 5 out of 10 classes. Compared to DeepInteraction, our approach improves construction worker by 25.2 AP, stroller by 6.8 AP, and pushable-pullable by 27.3 AP. Note, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and Stro. is stroller. We highlight classes with Medium and Few examples per class in blue.
Method Car Adult Truck CV Bicycle MC Child CW Stroller PP
@@
MMLF(CP, DINO) 86.3 86.2 60.6 35.3 70.1 75.9 8.8 55.9 37.7 58.1
TABLE XI: Evaluating 2D detection performance. We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes. However, 2D RGB detectors achieve higher performance. Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane. This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors. All detectors in this table are only trained on the nuScenes train-set.
@@
  1. (a)
    Ours (e.g., Finegrain score only)
  2. (b)
    Object score * Finegrain score ([87], e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$)
  3. (c)
    Coarse score * Finegrain score (Variant-1 of [87], e.g. $P_{CAR}=P_{VEHICLE}*P_{CAR}$)
  4. (d)
    Object score * Coarse score * Finegrain score (Variant-2 of [87], e.g. $P_{CAR}=P_{OBJ}*P_{VEHICLE}*P_{CAR}$)
Unlike [87, 88] which require a strict label hierarchy, our approach is not limited to a hierarchy. We find that other hierarchical methods improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
G Per-Class Breakdown Results
We compare recent multi-modal methods w.r.t per-class AP in Table X. All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community. Notably, our late-fusion approach achieves 20.0 higher AP on pushable-pullable and 6.0 higher AP on stroller than prior work. In general, our late-fusion approach yields considerable improvement on classes with medium and few examples. Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8 mAP, which is still much better than compared methods. We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
TABLE XII: AP vs. NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
FCOS3D BEVFormer CP TransFusion DeepInteraction CMT MMLF(CP, DINO)
[13] [23] [12] [34] [59] [10] (Ours)
mAP 20.9 27.3 39.2 39.8 43.7 44.4 51.4
NDS 30.4 38.8 54.9 53.9 54.4 55.9 60.4
TABLE XIII: 3D detection performance on the standard nuScenes benchmark. We evaluate the impact of late-fusion for the standard nuScenes benchmark. Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark. Notably, fusing CenterPoint [12] with DINO [11] improves detection accuracy by 2.6 mAP and 6.9 NDS. However, fusing CMT [10] (a multi-modal 3D detector) with DINO [11] increases detection accuracy by 0.2 mAP, only improving on bicycle and traffic-cone. We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion. Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
@@ G Per-Class Breakdown Results
MMLF(CMT, DINO) 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
H Analysis of 2D Detection Performance
Table XI shows that 2D detectors (e.g., DINO) outperform state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set. Importantly, DINO performs significantly better than BEVFormer (15.9 vs. 2.1 mAP) on rare classes.
I Analysis of AP vs. NDS Results
We report nuScenes Detection Score (NDS) and mAP results in Table XII. We find that all prior methods follow the same ranking on both metrics. This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
J Standard nuScenes Benchmark Results
We evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [6]. Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [12] and DINO [11] yields a 2.6 and 6.9 performance increase for mAP and NDS on the standard nuScenes benchmark, respectively. Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories. Next We apply our late-fusion to CMT [10], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements. Late-fusion of CMT [10] and DINO [11] only marginally improves bicycle and traffic cone performance. This suggests that CMT is already able to effectively utilize multi-modal information to correctly localize and classify common classes.
K State-of-the-Art Comparison on AV2
We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research. AV2 evaluates on 26 classes, which follow a long-tailed distribution. Following [9], we train on and evaluate detections up to 50m. As shown in Table XIV, our main conclusions from nuScenes still hold for AV2. FCOS3D [13] yields poor performance on all classes, likely due to inaccurate depth estimates. CenterPoint performs considerably better, achieving high accuracy on classes with many examples. Notably, CenterPoint performs better on AV2’s rare classes (30.2 mAP) compared to nuScenes’s rare classes (3.5 mAP), likely because AV2 has more examples per-class in-the-tail. Lastly, our proposed late-fusion approach yields a 8.3 mAPimprovement over CenterPoint and a 3.9 mAP improvement over prior work. These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Few       Medium      Many
Refer to caption Refer to caption
Figure 13: Left: According to the histogram of per-class object counts (on the left), classes in Argoverse 2.0 (AV2) follow a long tailed distribution. Following [1] and nuScenes, we report performance for three groups of classes based on their cardinality (split by dotted lines): Many, Medium, and Few. As AV2 does not provide a class hierarchy, we construct one by referring to the nuScenes hierarchy (cf. Fig. 13 on the right).
TABLE XIV: Comparison with the Argoverse 2 state-of-the-art. We present results on AV2 evaluated at 50m. FCOS3D [13] achieves poor performance, likely due to inaccurate depth estimates. In contrast, CenterPoint achieves strong performance on all classes. Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3 mAP improvement averaged over all classes. These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
@@ H Analysis of 2D Detection Performance
MMF(CenterPoint, FCOS3D) (Ours) C + L 48.4 79.0 51.4 35.3
MMLF(CenterPoint, DINO) (Ours) C + L 52.3 89.4 54.2 38.7